OpenStudy (rootbeer003):

In which direction does the parabola open? y = 3x^2 – 5x – 10 down up left right

TheSmartOne (thesmartone):

I'm here :)

OpenStudy (rootbeer003):

In which direction does the parabola open? y = 3x^2 – 5x – 10 down up left right

TheSmartOne (thesmartone):

look at the graph on here: https://www.wolframalpha.com/input/?i=y+%3D+3x%5E2+-+5x+-+10

OpenStudy (rootbeer003):

up

OpenStudy (anonymous):

x term is squared, leading coefficient is positive up is right

TheSmartOne (thesmartone):

bingo

OpenStudy (anonymous):

bingo

OpenStudy (rootbeer003):

Use the discriminant to determine how many x-intercepts the graph of the equation has. y = – 4x^2 + 3x + 2 zero one two three

OpenStudy (rootbeer003):

2?

TheSmartOne (thesmartone):

bingo

OpenStudy (anonymous):

\[b^2-4ac=3^2-4\times (-4)\times 2\] is positive so yeah two

OpenStudy (rootbeer003):

Where is the vertex of the parabola? y = x^2 + 2x + 3 On the y-axis On the x-axis Below the x-axis Above the x-axis

OpenStudy (rootbeer003):

d?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y+%3D+x^2+%2B+2x+%2B+3 what a nice picture

OpenStudy (anonymous):

yupo

TheSmartOne (thesmartone):

bingo

OpenStudy (anonymous):

@TheSmartOne cheated by writing "vertex"!!

TheSmartOne (thesmartone):

hehe

TheSmartOne (thesmartone):

just using wolfram is cheating ;)

OpenStudy (rootbeer003):

i aint

OpenStudy (rootbeer003):

jesus is whispering the answers

OpenStudy (rootbeer003):

Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m^2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden? 17 m 19 m 18 m 16 m

OpenStudy (anonymous):

name of my new country album

OpenStudy (anonymous):

i hate word problems

OpenStudy (rootbeer003):

nice

TheSmartOne (thesmartone):

one of the solutions is 3 x + 14 3 + 14 = ??

OpenStudy (rootbeer003):

17

OpenStudy (rootbeer003):

Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m^2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden? 17 m wide 16 m wide 14 m wide 15 m wide

OpenStudy (anonymous):

\[12+3=?\]

OpenStudy (rootbeer003):

ok next

TheSmartOne (thesmartone):

^

OpenStudy (anonymous):

same question but they wanted the width

OpenStudy (anonymous):

holly lol holly came from miami FLA

OpenStudy (rootbeer003):

Which equation would best help solve the following problem? Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m^2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? 255 = (12 + x^2)(14) 255 = (12)(14) 255 = (12 – x)(14 – x) 255 = (12 + x)(14 + x)

OpenStudy (anonymous):

deja vu that certain feeling you get when you think "i've been here before"

OpenStudy (anonymous):

deja vu that certain feeling you get when you think "i've been here before"

OpenStudy (anonymous):

deja vu that certain feeling you get when you think "i've been here before"

TheSmartOne (thesmartone):

the answer is in the link ttps://www.wolframalpha.com/input/?i=(x%2B12)(x%2B14)+%3D+255 that's the link I gave you 5 minutes ago

OpenStudy (anonymous):

same as link

OpenStudy (rootbeer003):

d?

OpenStudy (anonymous):

bimgo

TheSmartOne (thesmartone):

bingo* ;)

OpenStudy (anonymous):

bigmo

OpenStudy (anonymous):

where else can you get such great answers?

OpenStudy (anonymous):

er... help, i meant "help:

OpenStudy (rootbeer003):

from jesus

OpenStudy (rootbeer003):

The height of a triangle is 6 m more than its base. The area of the triangle is 56 m^2. Find the length of the base. 7 m 8 m 14 m 15 m

OpenStudy (rootbeer003):

8?

OpenStudy (anonymous):

idk let jesus answer

OpenStudy (rootbeer003):

its 8

TheSmartOne (thesmartone):

LOL

OpenStudy (rootbeer003):

is it 8

TheSmartOne (thesmartone):

https://www.wolframalpha.com/input/?i=0.5x(x%2B6)+%3D+56

TheSmartOne (thesmartone):

bingo

OpenStudy (rootbeer003):

ok im done

OpenStudy (anonymous):

are you sure about 8?

OpenStudy (anonymous):

8, 14, oh yeah

OpenStudy (rootbeer003):

im done time for criminology

TheSmartOne (thesmartone):

8 and -14

TheSmartOne (thesmartone):

bye

OpenStudy (anonymous):

enjoy your criminology

OpenStudy (anonymous):

and your sliders

TheSmartOne (thesmartone):

xD

OpenStudy (rootbeer003):

SHUEURHEUHREHRERU

OpenStudy (rootbeer003):

bye people

TheSmartOne (thesmartone):

cya

OpenStudy (anonymous):

ttyl as the kids say